OpenStudy (abbycross167):

The length of the hypotenuse of a right triangle is 15 ft. The length of one leg is 9 ft. Find the length of the other leg. A)6 B)21 C)12 D)18 My Answer ~ A or C

OpenStudy (shamim):

Do u know Pythagorean theorem

OpenStudy (abbycross167):

yes sir

OpenStudy (shamim):

Hypotenuse ^2=base^2+perpendicular ^2

OpenStudy (shamim):

15^2=9^2+leg^2

OpenStudy (abbycross167):

yes sir, I did that but I wasn't sure if I did the math right so i came here

OpenStudy (shamim):

Leg=?

OpenStudy (abbycross167):

225?

OpenStudy (shamim):

15^2-9^2=?

OpenStudy (abbycross167):

144?

OpenStudy (solomonzelman):

\(\large\color{#993300}{ \displaystyle a^2+b^2=c^2 }\) where \(\large\color{#993300}{ \displaystyle c }\) is the hypotenuse.

OpenStudy (solomonzelman):

You know one side is equal to 9 and the hypotenuse is equal to 15. Therefore, \(\large\color{#993300}{ \displaystyle a^2+9^2=15^2 }\)

OpenStudy (solomonzelman):

\(\large\color{#115599}{ \displaystyle a^2+9^2=15^2 }\) \(\large\color{#115599}{ \displaystyle a^2=15^2-9^2 }\) \(\large\color{#115599}{ \displaystyle a^2=225-81=144 }\) \(\large\color{#115599}{ \displaystyle a=\sqrt{144} }\)

OpenStudy (solomonzelman):

=12

OpenStudy (abbycross167):

Ok Thankyou kindly!!!

OpenStudy (solomonzelman):

yw

OpenStudy (phi):

you might want to memorize 3,4,5 are the legs of a right triangle (these numbers are used alot) you might also see 5,12,13 and, any multiple of 3,4,5 are also legs of a right triangle notice if we multiply each by 3 we get 9,12,15 this is the triangle in your problem